### v1

# Estimating the Influence of Sequentially Correlated Literary Properties in Textual Classification: A Data-Centric Hypothesis-Testing Approach

Abstract
Stylometry aims to distinguish authors by analyzing literary traits assumed to reflect semi-conscious choices distinct from elements like genre or theme. However, these components often overlap, complicating text classification based solely on feature distributions. While some literary properties, such as thematic content, are likely to manifest as correlations between adjacent text units, others, like authorial style, may be independent thereof. We introduce a hypothesis-testing approach to evaluate the influence of sequentially correlated literary properties on text classification, aiming to determine when these correlations drive classification. Using a multivariate binary distribution, our method models sequential correlations between text units as a stochastic process, assessing the likelihood of clustering across varying adjacency scales. This enables us to examine whether classification is dominated by sequentially correlated properties or remains independent. In experiments on a diverse English prose corpus, our analysis integrates traditional and neural embeddings within supervised and unsupervised frameworks. Results demonstrate that our approach effectively identifies when textual classification is not primarily influenced by sequentially correlated literary properties, particularly in cases where texts differ in authorial style or genre rather than by a single author within a similar genre.
1 Introduction
When considering the process of text composition, numerous properties come into play, with genre, style, and theme being particularly significant. Identifying the primary factor distinguishing texts among these three is crucial for tasks like authorship attribution and forensic linguistics (Holmes, 1994; Koppel and Schler, 2004).
Style, theme, and genre are fundamental properties that play distinct yet intertwined roles in the composition and interpretation of textual content. Style refers to the distinctive manner in which a writer expresses ideas, characterized by choices in vocabulary, sentence structure, tone, and rhetorical devices. It encompasses the unique linguistic fingerprint of an author and contributes to the overall aesthetic and readability of the text (e.g., Saukkonen, 2003). The theme, conversely, pertains to the central ideas, messages, or subjects explored within the text. It reflects the underlying concepts or motifs that recur throughout the narrative, conveying deeper meaning and resonance to the reader (e.g., Brinker, 1995). Genre, meanwhile, categorizes texts into distinct literary or textual forms based on shared conventions, structures, and expectations (e.g., Chandler, 1997). It provides a framework for understanding and classifying texts according to their overarching narrative structures, plot elements, and stylistic conventions.
In computational linguistics, characterizing specific desired properties of a text – such as theme, style, or genre – poses significant challenges due to their abstract and multifaceted nature Rybicki et al. (2016). These parameters encompass a broad range of linguistic elements, making them difficult to quantify and model computationally. Furthermore, these parameters vary widely across authors and texts, posing challenges for developing universal computational models (Serrano et al., 2009). Identifying patterns and connections within these parameters involves deciphering complex semantic and conceptual relationships, which can be context-dependent and subject to individual interpretation. This variability makes it challenging to automate analysis reliably across diverse texts. Additionally, the classification of texts based on these parameters faces similar challenges, as texts often defy traditional categorization and exhibit hybrid or evolving conventions. Additionally, such classification relies on intricate intertextual and extratextual factors, such as cultural context and reader expectations, which are challenging to formalize computationally. Overall, literary works’ abstract and multidimensional nature presents formidable obstacles in computational linguistics, demanding innovative approaches and interdisciplinary collaborations for adequate characterization.
Of particular interest is the discipline of stylometry, which presupposes the possibility of distinguishing between distinct authorial styles (that are independent of thematic content) based on statistical learning analyses Kestemont (2014). Identifying such a stylistic property is helpful in a variety of real-world settings (e.g., forensic linguistics Lambers and Veenman (2009), plagiarism detection (Ainsworth and Juola, 2018), and authorship analysis (Kestemont et al., 2016) – where it has been shown that other textual properties are often misleading in asserting authorship (Stamatatos, 2009, 2018)).
As with many statistical-learning tasks, approaches in stylometry are usually categorized into supervised and unsupervised settings Holmes (1994, 1998); Juola (2006). In the supervised case, several distinct authors and their works are provided (for training), and the task is to attribute texts to one of the authors under consideration, a task known as authorship attribution. Within this context, two distinct assumptions can be made: (1) that the author of each unattributed text is necessarily found within the given set of authors Koppel et al. (2009), and (2) that the composer of an unattributed text need not be among the authors under consideration Koppel et al. (2011). Alternatively, it may be desirable, given a single known author and some prior texts thereof, to verify whether some unknown work was also composed by this author or, instead, by a different one – a task known as authorship verification (e.g., Koppel et al., 2007; Juola, 2006; Kestemont et al., 2016). Other supervised settings include authorship profiling, where the labeling required is not for a distinct author but for a group with some shared profile (e.g., historical era, native language, gender, socio-economic class) Argamon et al. (2009). Furthermore, variants of these tasks include cross-genre settings, where the train set is from a different genre than the test set (e.g., prose-poetry Kestemont et al. (2012)), cross-domain settings (e.g., Facebook-Twitter Markov et al. (2018)), or both (Stamatatos, 2018). Lastly, there exists a variety of different case studies that incorporate different languages, historical eras, text sizes, and more (Schwartz et al., 2013).
In the unsupervised setting, no prior knowledge of the authors is provided. For a corpus of unlabeled texts (or a single text split into text units, such as sentences, paragraphs, or some arbitrary partition), it is required to estimate which texts were composed by the same author. At the root of this unsupervised task is the fundamental question: Given two texts, did the same author write them or not?
As the challenge at hand is identifying a typical style regardless of thematic content, literary features used in stylometry focus on the abundant rather than the scarce (Koppel and Winter, 2014). Unique standing-out words that allude to content are usually ignored, and preferred features are traditionally subsets of the most frequent words and sequences thereof, as these could attest to some authorial style that persists through different works regardless of content (e.g., Holmes, 1994; Argamon et al., 2007; Stamatatos, 2009; Escalante et al., 2011). Even today, when an interpretable analysis is of importance, commonly used features are $n$-gram features of words, characters ($k$-mers), part-of-speech tags (POS), or dependency parsing tags, which are usually applied with a frequency-dependent vectorization such as tf-idf Sammut and Webb (2010).
While some attention has been devoted to a thorough assessment of the utility of different literary features in various settings, most of this attention has focused on variants of the supervised setting of stylometry Koppel et al. (2013); Eder (2015); Juola et al. (2019). In the unsupervised setting, much less has been studied about the effects of different linguistic features and their dependence on textual properties such as genre or language-specificity. Moreover, the focus has been on assessing the contribution of different features and methods in achieving benchmarks. However, to the best of our knowledge, no prior work deals with ascertaining what part of the signal found can verifiably be attributed to different constituents of the literary signal, which can lead to false assertions concerning the results of such analyses.
Ensuring that an authorship attribution algorithm remains agnostic of qualities of the text that may be independent of style, such as theme or genre, is crucial for robust and accurate analysis (e.g., Eder, 2015; Rudman, 2016; Murauer and Specht, 2021). Thus, the algorithm can focus solely on identifying each author’s unique stylistic and linguistic patterns. This approach guards against biases arising from thematic similarities or differences between texts, which could obscure genuine authorial signatures. However, achieving such agnosticism presents challenges. Texts encompass a multitude of features, and separating the relevant stylistic cues from the thematic or genre-related elements – that are often closely coupled to style, requires sophisticated algorithms capable of discerning subtle linguistic nuances (e.g., Halvani et al., 2020). Moreover, a supervised algorithm must be trained on diverse datasets spanning various themes and genres to ensure its adaptability and effectiveness across different textual domains. Furthermore, employing interpretable methods is essential to ensure that the algorithm performs according to expectations, allowing researchers to understand and validate the reasoning behind its attribution decisions (e.g., Theophilo et al., 2022). Thus, balancing the need for agnosticism of non-stylistic content with the complexity of textual data poses a significant obstacle in developing authorship attribution algorithms that are both reliable and versatile.
In this work, we present and explore a novel approach that takes a first step towards identifying which constituent(s) of the literary signal (e.g., thematic, stylometric, genre-wise) is responsible for a distinction between a pair of texts. Specifically, we wish to rule out that the distinction is based on constituents that exhibit correlations between adjacent text units that include but are not limited to thematic content. Our method employs a hypothesis-testing framework that analyzes sequential correlations of label sequences of text units. By developing a stochastic model for these correlations, we provide an empirical basis for assessing whether sequentially correlated literary properties dominate classification outcomes. We demonstrate that relying solely on embeddings intended to capture stylistic features is often insufficient to rule out non-stylistic influences, especially when text units are short. Instead, our approach models sequential correlation distributions independently of any specific embedding or algorithm. Since adjacent text units tend to be correlated by, for example, thematic content, rather than necessarily by authorial style (in the event of distinct authors), they naturally exhibit similarity when embedded.
We consider pairs of texts made of text units of English prose of different similarity in genre, composed by either other or the same author. We conduct a lateral study by applying unsupervised and golden-standard supervised methods to distinguish between pairs of texts embedded using traditional and state-of-the-art neural embeddings optimized for authorship attribution. In each case, we apply our hypothesis-testing framework to determine whether the resulting classification is predominantly affected by sequential correlations – as we would expect for two texts written by the same author and are similar in genre and are therefore likely to cluster based on thematic distinctions – or not.
2 The State of Stylometry
Stylometry can be traced to the mid-20th century, when scholars such as Mosteller and Wallace employed statistical methods to analyze the writing styles of the Federalist Papers, demonstrating the effectiveness of word frequency and function word usage in distinguishing between authors (Mosteller and Wallace, 1963). Their pioneering work set the stage for applying machine learning and natural language processing techniques in stylometric research.
Advancements in computational linguistics and data science in the late 20th and early 21st centuries increased the scope of stylometric analysis expanded beyond authorship attribution to include authorship verification, genre classification, and text clustering (e.g., Juola et al., 2006; Juola, 2006; Rybicki and Heydel, 2013).
One of the landmark studies in contemporary stylometry by Koppel et al. (2007) explored the stylistic fingerprints of authors across different languages and genres. Using machine learning algorithms and linguistic features, they demonstrated the robustness of stylometric techniques in identifying individual writing styles and distinguishing between authors with high accuracy. There, they demonstrated and emphasized the importance of frequent features, such as common words or phrases, punctuation patterns (Darmon et al., 2021), and syntactic structures, in capturing the unique writing style of an author. They argued that these frequent features are more indicative of an author’s stylistic tendencies than rare or unique words, which may be more subject to variation or noise. Later works underpinned the fact that frequent features alone may not capture style comprehensively (e.g., Savoy, 2013; Lagutina et al., 2019; Grieve, 2023), and highlighted the importance of considering the limitations of frequency-based stylometric features and the need for complementary approaches to capture the richness and complexity of writing style.
Complementary methods to capture writing style beyond frequent features have been proposed to address the limitations of frequency-based stylometric analysis. One such approach involves incorporating syntactic and structural features of text, which can provide deeper insights into an author’s unique stylistic tendencies. Further studies have demonstrated the effectiveness of syntactic features, such as parse tree structures (e.g., Feng et al., 2012) and syntactic dependencies (e.g., Hollingsworth, 2012), in distinguishing between authors and genres. Additionally, semantic analysis techniques, which focus on the meaning and context of words and phrases, have been proposed as valuable complements to frequency-based methods. In conclusion, by integrating these complementary methods, researchers can achieve more comprehensive and accurate analyses of writing style in various textual domains.
More recently, stylometry experienced a significant evolution with the advent of Large Language Models (LLMs) and neural approaches (e.g., Ding et al., 2017; Canbay et al., 2020). Leveraging the power of large-scale pre-trained models such as the Generative Pre-trained Transformer (GPT) series, researchers have explored novel avenues in authorship attribution, genre classification, and other stylometric tasks (e.g., Huertas-Tato et al., 2024). Recent works showcase the capabilities of LLMs when applied to the authorship attribution task (e.g., Kumarage and Liu, 2023). However, despite their impressive performance in a wide variety of disciplines, as well as within the domain of language processing, many of these neural approaches lack interpretability, making it challenging to understand the underlying factors driving their predictions (Dror et al., 2018, 2020; Moulaison-Sandy, 2023). Moreover, in contexts such as historical texts where training data may be scarce or non-representative, these sophisticated models may not be applicable, highlighting the need for interpretable and domain-adaptive stylometric methods.
Despite the rapid and impressive growth in textual analysis for a myriad of purposes, there has been a noticeable scarcity of works specifically aimed at discerning various literary parameters, including theme, style, and genre, to ascertain which one predominantly drives text classification (Lagutina et al., 2019; Litvinova, 2020; Hou and Huang, 2020). This lack of focus on disentangling the distinct influences of these parameters – which is difficult and often impossible due to the inherent vaguely defined characteristics thereof – can lead to potential false positive classifications based solely on a hypothesized factor (e.g., Schuster et al., 2020). Therefore, while stylometric analysis traditionally prioritized tasks such as authorship attribution, a broader consideration of literary parameters is essential for accurate text classification and a deeper understanding of textual characteristics.
3 Hypothesis-Testing Scheme
In the context of textual data analysis, sequential correlations are believed to manifest at different adjacency scales due to the inherent structural organization of the text. Textual content often exhibits sequential coherence and organization, where related concepts, ideas, or themes are often clustered together within contiguous segments of text (e.g., Ghadessy and Gao, 2001; Kim and Wilbur, 2012; Armborst, 2017). At a finer adjacency scale, adjacent words or phrases within a sentence or paragraph are more likely to be sequentially related, capturing local textual coherence. On larger adjacency scales, such as between sentences, paragraphs, or chapters, thematic correlations may evolve to capture broader sequential shifts or transitions within the text. For instance, transitions between paragraphs may signal shifts in topic or theme, while transitions between chapters may denote major shifts in narrative focus or subject matter. By considering sequential correlations at these different adjacency scales, we can gain a more comprehensive understanding of the text’s underlying sequential structure and organization, enabling more accurate analysis and interpretation of its classification.
Recently, we suggested in Yoffe et al. (2023) that correlations between adjacent text units affect their classification based on thematic similarity, potentially masking stylistic differences. Consequently, when devising a hypothesis-testing framework to estimate the statistical significance of a potentially stylistic distinction between texts, the assumption of naive label permutation, which ignores these correlations and assumes label independence, is unsuitable for textual data. We proposed that human experts’ labeling tendencies may reflect a predisposition to assign authorship to text units sharing sequential similarities, as often observed in Biblical exegetical analysis (e.g., Dershowitz et al., 2015). This suggests that when human experts annotate text segments, they tend to cluster together words, phrases, or passages with thematic coherence, interpreting them as originating from the same authorial source. There, we made the statistical argument that shifting the labels cyclically conserves predominantly sequential correlations. Indeed, this cyclic-shift test is a parochial solution and is often unsuitable for similar tasks since even if a human expert’s labeling is available – its underlying rationale can be affected by a convoluted mixture of literary parameters.
Here, we propose a data-centric approach that assumes labels to be correlated random variables in the sequential domain (what we refer to as label space). By only being concerned with label sequences, this approach is essentially embedding-invariant and can be applied to texts (or other datasets) embedded densely, sparsely, or with any desired subset of features or neural or traditional language model. By estimating the intrinsic average correlations between text units at different lags (i.e., distances along the textual sequence), we generate a multivariate distribution from which we draw label sequences that are accordingly correlated between text units at different sequential scales. This approach enables us to sample a random set of labels in a way that simulates the inherent sequential correlations apparent in texts and, therefore, provides a null hypothesis with such characteristics.
In this work, we consider the null hypothesis labeling-generation routine for the simplest scenario of a text composed by one or two authors. Once this sampling mechanism for our null distribution is established, we can compute the $p$-value indicating the likelihood that the distinction between the two parts of the text indicated by the labels of $L$ is made due to classification driven by sequentially correlated literary properties.
Reference label sequence $L^{(\rm{real})}$ of a textual data set
We consider correlations in the label space. Let $D\in\mathbb{R}^{m\times f}$ denote a textual data set containing $m$ samples (text units – a sequence of words of some length) of $f$ dimensions (features). We denote a sequence of labels representing some hypothesized or real partition between texts (or an internal partition within a single text, as discussed in Yoffe et al. (2023)) as $L^{(\rm{real})}$.
Label sequence $L$ of a classification algorithm applied to $D$
We apply some classification algorithm with two labels to $D$ and receive a sequence of labels $L\in\{0,1\}^{m}$.
Auto-covariance of $L$
We proceed to compute the average auto-covariance of $L$ for all lags, to receive a vector $A\in\mathbb{R}^{m}$, where $A_{\ell}$ ($\ell=1,\ldots,m$) is the average auto-covariance of $L$ at lag $\ell$, given by
$A_{\ell}=\frac{1}{m}\sum_{j=1}^{m-\ell}(L_{j}-\bar{L})(L_{j+\ell}-\bar{L}),$ (1)
where $\bar{L}=1/m\sum_{i}L_{i}$ is the average of the elements of $L=(L_{1},\ldots,L_{m})^{T}$. To impose linear decay as a function of lag (i.e., that the strength of the correlation between farther separated text units decays linearly), all terms in $A$ are scaled by $1/m$, rather than the expected term for averages $1/(m-\ell)$. In other words, Eq. (1) is equivalent to
$A_{\ell}=\frac{m-\ell}{m}\left[\frac{1}{m-\ell}\sum_{j=1}^{m-\ell}(L_{j}-\bar{L})(L_{j+\ell}-\bar{L})\right],$
which approaches zero linearly as $\ell\to m$.
Average auto-covariance matrix $M$
Let $M\in\mathbb{R}^{m\times m}$ denote the average autocovariance matrix, whose entries are defined by
$(M)_{ij}=A_{|i-j|}.$ (2)
Thus, the correlation between every $i$th and $j$th entries in $L$ is given by the (linearly decaying) average autocovariance of $L$ at lag $|i-j|$. This approach represents the empirical expectation value of the global correlation trend in the text of every two labels at some fixed lag of size $|i-j|$.
Intuition: The average autocovariance matrix $M$ serves as a structured representation of the typical correlations (i.e., $A_{\ell}$) between labels at varying sequential distances ($\ell$) within the dataset. The $i$th row in $M$ corresponds to the linearly-decaying autocovariance of the $i$th label ($(M)_{ii}$) with respect to all other labels in the dataset ($(M)_{ij\neq i}$), which is governed by (1): their proximity to it (i.e., lag), and (2) the average autocovariance between all labels in the dataset separated by that lag. This means that entries closer to $(M)_{ii}$ represent correlations between the $i$th label and its closer neighbors, and they would therefore be more correlated thereto than farther entries. Thus, $M$ specifically captures the average autocovariance between labels at different sequential distances within the dataset – the necessary mechanism to model sequential correlations. As opposed to $M$, a simple covariance matrix captures the covariance between pairs of labels in the dataset without considering their sequential arrangement. Each entry in the covariance matrix quantifies pairwise relationships between every two labels, irrespective of their positions within the sequence.
Generating a correlated label sequence $L^{(null)}$
We use the autocovariance matrix $M$ to draw a random sequence of correlated labels $L^{(\rm{null})}\in\mathbb{R}^{m}$ in the following manner: consider the multivariate normal distribution $V\sim\mathcal{N}(\overrightarrow{\bar{L}},M)$, where $\overrightarrow{\bar{L}}=(\bar{L},\ldots,\bar{L})^{T}\in\mathbb{R}^{m}$. The choice of $V$ as a multivariate normal distribution stems from the ease of embedding the average autocovariance matrix $M$ as a covariance matrix. Let $X=(X_{1},...,X_{m})\in\mathbb{R}^{m}$ be drawn from $V$, which can be transformed into a vector of correlated labels ($L^{(\rm{null})}$) using elementwise thresholding (Leisch et al., 1998):
$L_{i}^{(\rm{null})}=\begin{cases}1,&X_{i}>0.5\\ 0,&\textrm{otherwise}.\end{cases}$
Essentially, this means that the probability $p_{i}$ of the $i$th label in $L^{(\rm{null})}$ to be 1 is given by
$p_{i}\equiv\mathbb{P}(L^{(\rm{null})}_{i}=1)=\mathbb{P}(X_{i}>0.5).$ (3)
Intuition
The multivariate normal distribution $V$ is a convenient method to couple the two elements that are essential for generating stochastic correlated label sequences: (1) the probability of each label to be 1, which is identical for all entries and is simply the average label value of the reference label sequence $L$, and (2) the average autocovariance matrix $M$, determining how correlated each entry in the vector $X$ that is drawn from $V$ is with all other labels. $X$, in turn, is a specific realization drawn from $V$, whose each entry is a continuous number, affected by correlations with the entire dataset. Lastly, elementwise thresholding ensures that the values of $X$ assume the binary form of a label sequence – either 0 or 1.
Thus, $V$ enables the generation of stochastic label sequences that preserve sequential dependencies that exist in the reference label sequence $L$, thereby facilitating a robust method to generate a null distribution that would examine the measure of agreement between sequentially correlated label sequences and a real or hypothesized partition between two texts.
Within a hypothesis-testing framework, we may probe the null hypothesis whereby the reference label sequence $L$ is one such realization of labels that only exhibit sequential correlations. Thus, assuming that $L$ is characterized primarily by sequential correlations, we use it to generate the underlying distribution from which it was presumably sampled. Using this distribution, we generate stochastic label sequences that possess similar sequential qualities. In the event that the agreement between the real label sequence $L^{(\rm{real})}$ reference label sequence $L$ is not based on sequential correlations (but rather on, e.g., stylistic differences) – that is, that the null hypothesis is wrong, the resulting overlap between the null label sequences and $L^{(\rm{real})}$ will not be comparable to its overlap with $L$ since the bulk of agreement between the two originates from a non-sequential axis. This comparison is insightful regarding the extent to which sequential correlations contribute to the observed label assignments, underpinning the robustness and significance of non-sequentially correlated distinctions between texts. In Figure 1, we provide a visualization of $A$ and $M$ for some label sequence $L$ of length 1300 to provide some additional intuition.
Figure 1: Vizualization of the hypothesis-testing framework’s building-blocks: Left Panel: Average auto-covariance vector $A$ computed for some label sequence of size 1300. The $i$th entry is $A_{\ell=i}$ (see Eq. 1). Right Panel: Average auto-covariance matrix $M$ (see Eq. 2) The insert visualizes the 500th row in $M$, which is a mirror image of $A$ around the 500th entry, mapping correlations between that label and all other labels in the sequence.
Matthews Correlation Coefficient (MCC)
We consider the MCC model evaluation metric, given by
$MCC=\frac{TN\times TP-FN\times FP}{\sqrt{(TP+FP)(TP+FN)(FN+FP)(TN+FN)}},$ (4)
where T (F) stands for True (False) and P (N) stands for Positive (Negative), which are evaluations of some sequence against another. We use MCC to evaluate the agreement between the reference label sequence $L^{(\rm{real})}$ and $L$, or a drawn label sequence $L^{(null)}$ as our statistic of choice, normalized to percent within the range of 50%–100%, where 50% suggests an arbitrary overlap between two label sequences, and 100% suggests perfect overlap. By drawing multiple random correlated label sequences $L^{(\rm{null})}$, we empirically estimate the null distribution of MCC between the random and $L^{(\rm{real})}$. We can derive the probability of the MCC score between the latter and $L$ to be drawn from that distribution – implying that the unsupervised partition is based predominantly on sequentially correlated literary properties.
We present the procedure of generating a null label sequence $L^{\rm{(null)}}$ using this method schematically in Algorithm 1.
Algorithm 1 Hypothesis-testing Scheme: Generating a null label sequence $L^{\rm{(null)}}$
1:Apply a classification algorithm on an embedded corpus $D\in\mathbb{R}^{m\times f}$ and receive a label sequence $L\in\mathbb{R}^{m}$.
2:Calculate the (decaying) average autocovariance array $A$ of $L$, where $A_{\ell}$ is computed with Eq. 1, and $\ell\in\{1,\ldots,m\}$.
3:Generate the average autocovariance matrix $M$, such that $(M)_{ij}=A_{|i-j|}$.
4:Generate a multivariate normal distribution $V\sim\mathcal{N}(\overrightarrow{\bar{L}},M)$, where $\overrightarrow{\bar{L}}=(\bar{L},\ldots,\bar{L})\in\mathbb{R}^{m}$.
5:Draw a vector $X=\{X_{1},\ldots,X_{m}\}$ from $V$.
6:Impose elementwise thresholding on $X$, such that $X_{i}=1\Leftrightarrow X_{i}>0.5$, and $X_{i}=0$ otherwise.
7:$X$ is now $L^{\rm{\rm{(null)}}}$.
4 Experiments
4.1 Rationale
Our research is targeted primarily to assist in the analysis of historical texts, which are often multilayered and redacted. Unlike the analysis of modern texts (whose authors are known) predominantly used to explore the underpinnings of linguistic patterns and structure in literary works, studying ancient texts often serves a different purpose. Specifically, tracing the text’s origin to its extant version often requires interdisciplinary efforts across linguistics, history, and computer science disciplines (e.g., Bühler et al., 2024).
Given the nature of such endeavors, where insights from multiple fields converge, it is imperative to employ interpretable methods. Historical texts pose unique challenges that demand a heightened focus on interpretability in analysis (Chapman, 2017). They often contain layers of annotations, marginalia, and modifications made by scribes or editors over time, obscuring the original authorial intent (e.g., Müller et al., 2014). Additionally, historical texts may be written in archaic or obsolete languages, further complicating automated analysis (e.g., Baum, 2017). Moreover, the historical context surrounding these texts, including cultural norms, linguistic conventions, and socio-political factors, must be carefully considered to ensure accurate interpretation.
Given these complexities, employing interpretable methods is essential for unraveling the intricate layers of meaning embedded within historical texts (Chapman, 2017). This approach enables researchers to navigate the nuances of language and context for a deeper understanding of historical narratives and societal dynamics.
Therefore, rather than solely relying on state-of-the-art neural language models, we emphasize the exploration of ”older” features that have proven efficacy in textual analysis. By prioritizing interpretability and drawing on established methodologies, our approach ensures that the insights obtained from our study are accessible and meaningful across disciplinary boundaries. This commitment to methodological transparency and interpretability strengthens the interdisciplinary dialogue and enhances the applicability of our research findings.
That said, by leveraging traditional textual analysis techniques alongside modern neural embeddings, we aim to discern how well each method captures the unique stylistic elements. Through supervised experiments, we train models to differentiate between pairs of texts that were either composed by the same author or not. Specifically, we wish to explore potential biases inherent in supervised approaches, as they may rely on labeled data that could inadvertently reflect the perspectives or assumptions of annotators.
In contrast, unsupervised experiments allow for the exploration of the underlying structure of texts without predefined annotation. By applying clustering algorithms to traditional and neural embeddings, we can uncover patterns of stylistic variation and similarity within our dataset. This approach enables us to examine whether neural embeddings offer a more nuanced representation of stylistic features, thereby facilitating a deeper understanding of the authorship of texts.
Through these experiments, coupled with our hypothesis-testing approach, we seek to demonstrate the strengths and limitations of both traditional and neural methods in capturing the stylistic components of texts and the susceptibility of supervised approaches to training bias under different embedding setups.
Lastly, we aim to demonstrate that our hypothesis-testing approach provides critical insights into the basis of textual classification results, substantially enhancing interpretability. By quantifying the extent to which classification depends on sequentially versus non-sequentially correlated features, our method enables a nuanced understanding of the factors driving distinctions between texts. This deeper analysis clarifies when specific elements, such as thematic content, rather than authorial style, play a primary role, allowing researchers to discern the true drivers of classification outcomes.
4.2 Corpora
We perform our analysis on several corpora of English prose. These are (1) the Harry Potter series [^1] Freely available at https://github.com/formcept/whiteboard/tree/master/nbviewer/notebooks/data/harrypotter by J. K. Rowling, (2) four books of the Percy Jackson series by Rick Riordan, [^2] Freely available at https://archive.org/details/PercyJacksonTheLightningThief (3) works of J. R. R. Tolkien, including the The Lord of the Rings trilogy and The Silmarillion, [^3] Freely available at https://archive.org/details/TheSilmarillionIllustratedJ.R.R.TolkienTedNasmith/ (4) select works by Charles Dickens, [^4] Freely available at https://www.gutenberg.org (5) select works by C. S. Lewis, [^5] Freely available at https://gutenberg.ca/ebooks (6) select works by Jane Austen, [^6] Freely available at https://archive.org/ (7) select works by Arthur Conan Doyle, [^7] Freely available at https://sherlock-holm.es/ascii and (8) a ”spooky” corpus [^8] Freely available at https://www.kaggle.com/competitions/spooky-author-identification/data of fiction by Edgar Allan Poe, H. P. Lovecraft and Mary Shelley. We list all texts, their respective authors, and abbreviations used in this work in Table 1. All texts were preprocessed to contain only uncontracted words. [^9] Preprocessing code available in https://github.com/YoffeG/Thematic-nonThematic_Hypothesis_Testing Although punctuation was shown to bear significant stylistic merit (Lagutina et al., 2019; Darmon et al., 2021), we did not consider it in this work from the desire for this analysis to apply to texts that are not punctuated as well (see §4.1).
The chosen corpora allow for a comprehensive examination of the distinguishing power of our null hypothesis under different textual regimes of English prose, such as, for example, two distinct authors, distinct works of the same author of a similar/different genre (e.g., The Lord of the Rings series versus The Silmarillion (Upstone, 2002), the Narnia series versus The Screwtape Letters, intrinsic variance within Charles Dickens’ novels (Miller, 1958)), and (disjointly sampled) identical work of the same author. While not exhaustive, this sample may sufficiently represent different modes of interplay between theme, style, and genre and how they manifest in English prose text classification tasks – and, in turn, which modes can be discerned by our proposed method.
In this experiment, we attempted to distinguish 50 random pairs of works by different authors from all of the corpora above and all pairwise permutations (55) of works by the same author. Table 1 lists all texts used throughout this study, their authors, and their abbreviations as were used in this study.
The process for generating a sample of 50 pairs of texts, each authored by different individuals, permits flexibility in the representation of authors’ works. Initially, all works are grouped by their respective authors, except those inherently distinct, such as the The Lord of the Rings Trilogy and The Silmarillion. When selecting the two distinct authors for comparison, the process randomly determines the number of works to include in each author’s sample, with a maximum of three books per author, and also randomly selects which works constitute the sample. These author groups and their corresponding samples are listed in Appendix A. Sampling multiple books per author allows for a more comprehensive representation of each author’s stylistic range and tendencies. By considering a diverse selection of works from each author, we aim to capture a broader spectrum of their writing style, encompassing variations in themes, narrative structures, and linguistic nuances. This approach enhances the robustness of our comparative analysis, enabling a more nuanced understanding of authorial styles across different texts. Additionally, by including multiple books, we acknowledge the interconnectedness of themes and motifs within an author’s body of work, providing insights into recurring patterns and sequential evolution over time.
In Appendix A, we list two sets of data: the first containing 50 randomly selected pairs of text samples by different authors, and the second comprising all 55 permutations between texts authored by the same individual. In the latter table, we also denote our assumptions regarding the extent of genre variation within each pair of texts. When unsure, we mark such pairs with a ”?”. These analyses are conducted across four configurations, distinguishing between unsupervised and supervised approaches applied to traditionally and neurally-embedded texts, which are elaborated upon in subsequent sections.
Text Title Author Abbreviation
Harry Potter and the Philosopher’s Stone J. K. Rowling Potter_1
Harry Potter and the Chamber of Secrets J. K. Rowling Potter_2
Harry Potter and the Prisoner of Azkaban J. K. Rowling Potter_3
Harry Potter and the Goblet of Fire J. K. Rowling Potter_4
Harry Potter and the Order of the Phoenix J. K. Rowling Potter_5
Harry Potter and the Half-Blood Prince J. K. Rowling Potter_6
Harry Potter and the Deathly Hallows J. K. Rowling Potter_7
Percy Jackson and the Olympians: The Lightning Thief Rick Riordan Percy_1
Percy Jackson and the Olympians: The Sea of Monsters Rick Riordan Percy_2
Percy Jackson and the Olympians: The Titan’s Curse Rick Riordan Percy_3
Percy Jackson and the Olympians: The Demigod Files Rick Riordan Percy_5
The Lord of the Rings: The Fellowship of the Ring J. R. R. Tolkien Lotr_1
The Lord of the Rings: The Two Towers J. R. R. Tolkien Lotr_2
The Lord of the Rings: The Return of the King J. R. R. Tolkien Lotr_3
The Silmarillion J. R. R. Tolkien Silmarillion
David Copperfield Charles Dickens Dickens_Copperfield
Oliver Twist Charles Dickens Dickens_OliverTwist
Hard Times Charles Dickens Dickens_HardTimes
A Tale of Two Cities Charles Dickens Dickens_TwoCities
The Pickwick Papers Charles Dickens Dickens_Pickwick
The Adventures of Sherlock Holmes Sir Arthur Conan Doyle Holmes_1
The Memoirs of Sherlock Holmes Sir Arthur Conan Doyle Holmes_2
The Return of Sherlock Holmes Sir Arthur Conan Doyle Holmes_3
The Jungle Book Rudyard Kipling Kipling_JungleBook
The Second Jungle Book Rudyard Kipling Kipling_JungleBook2
Barrack-Room Ballads Rudyard Kipling Kipling_Ballads
The Chronicles of Narnia C. S. Lewis Narnia
Prince Caspian C. S. Lewis Caspian
The Screwtape Letters C. S. Lewis Screwtape
Emma Jane Austen Austen_Emma
Pride and Prejudice Jane Austen Austen_PP
Mansfield Park Jane Austen Austen_MansfieldPark
Collection of works (spooky corpus) Edgar Allen Poe spooky_EAP
Collection of works (spooky corpus) Marry Shelley spooky_MWS
Collection of works (spooky corpus) H. P. Lovecraft spooky_HPL
Table 1: All texts used in this work, their respective author, and abbreviation when referred to.
4.3 Preprocessing
In our preprocessing pipeline, we employ several techniques to standardize and prepare the text data for embedding. First, we convert all texts to lowercase to ensure uniformity and mitigate the impact of case sensitivity on subsequent processing steps. Additionally, we systematically remove punctuation marks to focus solely on the textual content. Following this, we concatenate all the texts into a single continuous string, facilitating the segmentation of the corpus into text units of a desired length.
In the traditional embedding setup (§4.4.1), all texts undergo named entity recognition and removal using spaCy’s entity extraction routine [^10] All entities apart from the following are removed: CARDINAL, ORDINAL, QUANTITY, PERCENT, TIME, DATE, LANGUAGE, PRODUCT. (Vasiliev, 2020), to decrease the number of features that are of predominantly thematic and semantic merit (Ríos-Toledo et al., 2017).
To simulate realistic document mixing, we split the second text into segments based on a Poisson distribution with $\lambda=3$, which introduces variability in the length of the segment. Then these segments are randomly inserted into various locations within the first text, rather than concatenated linearly. [^11] A list of how each pair of texts was commingled can be found at https://github.com/YoffeG/Thematic-nonThematic_Hypothesis_Testing This method enhances the realism of document integration by mimicking potential co-authored document interspersion. It avoids oversimplified models and tests the robustness of our approach under various conditions. A $\lambda$ value of 3 strikes a balance by avoiding excessively short segments that could disrupt coherence, while still allowing larger (and fewer) blocks to be effectively probed.
4.4 Embedding
4.4.1 Traditional Embedding Setup
In this setup, we embed the text using the traditional tf-idf vector-space representation (Aizawa, 2003) for each combination of these parameters. The tf-idf embedding technique has been demonstrated by works such as Fabien et al. (2020); Marcińczuk et al. (2021) to be appropriate, and often optimal, in tasks of unsupervised classification of texts, when neural language models are not implemented, either by choice or by necessity. In Yoffe et al. (2023), for example, we demonstrated how, by leveraging on the coupled properties of the $k$-means loss and tf-idf embedding, we could extract cluster-associated features that signify their importance to the classification, as well as their provenance as individual features or parts of a sequence. Each pair of texts is embedded according to a combination of several parameters across a grid of all permutations thereof: $n$, and $\ell$, that span across the following ranges: $n\in\{1,2,3,4\}$ for word $n$-grams and $n\in\{1,2,3,4,5,6\}$ for character $k$-mers, $\ell\in\{10,50,100,250,500,750,1000,2000\}$. This scope is motivated by the desire to exhaustively cover potential feature spaces while ensuring that the extracted features remain both meaningful and non-sparse (e.g., Allison et al., 2006; Antonia et al., 2014). By considering word $n$-grams and character $k$-mers within these ranges, we aim to capture a wide array of linguistic patterns and structural nuances present in the text data. Furthermore, this approach enables us to explore varying levels of granularity, from individual characters to multi-word phrases, thereby enriching the representation of the text corpus.
In this embedding setup, consider feature subsets consisting of the $f\in\{100,200,300\}$ most corpus-wise-weighed frequent joint words (i.e., the words whose average frequency across both corpora is the highest) while omitting the rest, generating $n$-grams from the post-omission text – and of those keeping the $f$ most frequent ones (i.e., $f$ most frequent $n$-grams generated from the text from which all but the $f$ most frequent words were omitted). The rationale behind selecting the top $f$ most corpus-wise-weighed frequent joint words and generating $n$-grams from the resultant text lies in the attempt to capture the most prominent linguistic features while minimizing noise for stylistic analysis (Akiva and Koppel, 2012). By computing the average frequency of words across both corpora and selecting the top $f$ words based on this criterion, we prioritize consistently prevalent words in both corpora, ensuring a robust representation of shared stylistic elements. This approach helps mitigate the potential bias towards words that are individually frequent in either corpus but may not necessarily reflect the stylistic characteristics common to both. Additionally, generating word $n$-grams/character $k$-mers from the pruned text further refines the analysis by capturing contextual dependencies and patterns characteristic of the selected corpus-wise-weighed frequent words. By retaining only the most frequent $n$-grams, we focus on capturing recurring linguistic structures indicative of the stylistic nuances in the text (Lagutina et al., 2019).
4.4.2 Neural Embedding Setup
We use an off-the-shelf implementation of the STAR [^12] Available at: https://huggingface.co/AIDA-UPM/star (Style Transformer for Authorship Representations) embedding (Huertas-Tato et al., 2023) to test the validity of our approach when using a state-of-the-art authorship-attribution-oriented neural embedding. STAR is a contrastively trained model fit to learn authorship embeddings instead of semantics. Similarly to the traditional embeddings, we implement STAR in two scenarios: (1) the unsupervised approach (§4.5.1), whereby we cluster texts using a $2$-means clustering algorithm on STAR-embedded pairs of texts and apply our hypothesis-testing framework to that classification. (2) a supervised cosine similarity approach (§4.5.2), as implemented in Huertas-Tato et al. (2023). In this case, a comparison between the traditional style-bearing features and STAR emphasizes the potential advantage novel neural models that are optimized to preserve and isolate the stylistic signal may have therein.
4.5 Classification
4.5.1 Unsupervised Approach
We use $k$-means as our clustering choice, with $k=2$ – a relatively standard and straightforward clustering algorithm that allows for helpful feature importance analysis (see §2.8 in Yoffe et al. (2023)).
Each text pair is evaluated with various word $n$-gram/character $k$-mer sizes and text unit lengths. The final embedded corpus is represented as $T^{(\ell,n,f)}$, where $n$, $\ell$, and $f$ denote the number of unique $n$-gram/$k$-mers, text unit length, and the number of features, respectively. The $2$-means algorithm is then applied to generate labels, and the MCC score quantifies the overlap between these labels and the original partition.
To ensure the statistical stability and robustness of the classification to different combinations of text units, we perform the procedure mentioned above by batch-sampling 20 subsamples of $c\cdot\rm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ randomly-drawn text units (for each of the two texts), where $c\in(0,1)$ is some factor of the size of the subsample in terms of the number of text units in the shorter text. The rationale behind this manner of batch-sampling (rather than, for example, sampling text units in sequences) procedure stems from the following empirical observation: Texts contain within them local variances in feature distributions (e.g., sequences containing mainly first/third person narration), that, when sampled sequentially, may yield a feature distribution that is unrepresentative of the global text-wise feature distribution, and thus misrepresent it. The batch-sample and subsample size factor ($c$) values were chosen by performing a statistical stability test of MCC distributions for several nominal parameter combinations and pairs of texts – and in this case, chosen to be $c=0.2$. Once a sub-sample is chosen, it is ordered according to the relative position of the text units in the text sequence (i.e., argsort) to retain sequential correlations at the arbitrarily chosen scales.
In every batch-sampling iteration, the subsample is submitted to the $2$-means, outputting the unsupervised label sequence $L$, whose MCC score is computed with respect to the actual partition $L^{\rm{(\rm{real})}}$. The $p$-value of this MCC score is then estimated with our hypothesis-testing routine (see §3), where we generate the null distribution of MCCs by drawing correlated label sequences $L^{\rm{(null)}}$ and compute their MCC score with respect to $L^{\rm{(\rm{real})}}$. In this study, we generate the null distribution of MCC scores out of 1000 sequences of $L^{\rm{(null)}}$ for each batch-sampling iteration.
Lastly, once the batch-sampling procedure is complete, we receive the empirical distribution of MCCs and $p$-values for that particular combination of parameters. The latter are corrected by the Benjamini-Hochberg False-discovery Rate (FDR) procedure for independent test statistics (Benjamini and Hochberg, 1995). Consequently, for every combination of $n$ and $\ell$, we compute the average MCC score and FDR-corrected $p$-value as representative values. The unsupervised experiment is summarized for some parameter combination of $n$ and $\ell$ and a feature space $f$ in Algorithm 2.
Algorithm 2 Experiment: Unsupervised ($k$-means)
1:Initiate MCC score array $v$ of length 100.
2:Initiate $p$-value array $u$ of length 100.
3:for $i=1,\ldots,100$ do
4:     Draw a subsample of $c\cdot\rm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units ($T^{(\ell,n,f)}_{\rm{subsample}}$).
5:     Apply $2$-means on $T^{(\ell,n,f)}_{\rm{subsample}}$ and receive label sequence $L$.
6:     Compute MCC score between $L$ and $L^{\rm{(\rm{real})}}$ and store in $v$.
7:     Estimate $p$-value of the MCC score with hypothesis-testing routine (null distribution composed of 1000 simulations) and store in $u$.
8:end for
9:Apply FDR correction to $u$ to receive an FDR-corrected array $u_{\rm{FDR}}$.
10:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\rm{FDR}}$.
4.5.2 Supervised Approach (General Imposters/Cosine Similarity Methods)
Here, our goal is to test the susceptibility of supervised classification methods to false-positive classifications of texts that are attributed to, for example, differences in authorial style, while being indeed affected by sequentially correlated literary properties, such as thematic content. This susceptibility arises because the chosen subsets of features often encapsulate multiple facets of the text, such as vocabulary, syntax, and punctuation, leading to a blending of stylistic and genre-wise characteristics within the classification process. Moreover, traditional features are not optimized to represent some desired literary property; rather, they are selected based on theoretical frameworks or limited empirical findings, potentially overlooking nuanced stylistic distinctions. We consider two supervised classification approaches: (1): The General Imposters [^13] We used the version available at https://github.com/bnagy/ruzicka, originally developed within the scope of the work of Kestemont et al. (2016). (GI) framework (Koppel and Winter, 2014) – a gold-standard supervised method used in multiple authorship-attribution scenarios (e.g., Kestemont et al., 2016; Juola, 2021). The GI model is designed to account for the possibility that authors might mimic or imitate the writing styles of others, either intentionally or unintentionally. This means the model doesn’t just rely on a single set of features to identify an author’s style but rather considers multiple feature subsets that capture various aspects of writing style. (2): Cosine similarity. In this approach, as implemented in Huertas-Tato et al. (2023), classification is performed by attributing an embedded text unit to either one or another centroid of a training-set sample of each class (using the original partition $L^{\rm{real}}$) by measuring the cosine similarity between them. In this case, all features are used to compute the cosine similarity.
Neural embeddings, unlike traditional features such as word $n$-grams, inherently capture diverse linguistic properties without manual selection. Attempting to under-sample or select subsets of features from neural embeddings, as in the GI framework, may not be suitable due to their holistic representation of text. Utilizing all features in neural embeddings is often more effective for classification, leveraging their rich linguistic representations encoded in high-dimensional spaces. Therefore, in our analysis, we apply the GI framework to classify the traditionally-embedded texts, whereas in the neurally-embedded case, we apply the cosine similarity classification. We set a consistent train/test split ratio of 20%–80% in all supervised scenarios to evaluate the model’s performance and assess the model’s generalization ability without favoring either the training or test.
The supervised classification procedure using the GI and cosine similarity frameworks for some parameter combination of $n$ (only in the GI case) and $\ell$ and a feature space $f$ is summarized in Algorithms 3 and 4, respectively.
Algorithm 3 Experiment: Supervised (GI)
1:Initiate MCC score array $v$ of length 100.
2:Initiate $p$-value array $u$ of length 100.
3:for $i=1,\ldots,100$ do
4:     Draw a subsample of $c\cdot\rm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units for training ($T^{(\ell,n,f)}_{\rm{train}}$), the rest are for testing ($T^{(\ell,n,f)}_{\rm{test}}$).
5:     Compute two label-wise centroids of $T^{(\ell,n,f)}_{\rm{train}}$ according to $L^{\rm{(\rm{real})}}$.
6:     Initiate label-association array $w\in\mathbb{R}^{100\times|T^{(\ell,n,f)}_{\rm{test}}|}$.
7:     for $j=1,\ldots,100$ do
8:         Draw a subsample of $f$, satisfying $f_{\rm{subsample}}\in f$ and $|f_{\rm{subsample}}|\approx 0.3\cdot|f|$.
9:         for $k=1,\ldots,|T^{(\ell,n,f)}_{\rm{test}}|$ do
10:              Compute the $2$-norm of the difference between $T^{(\ell,n,f_{\rm{subsample}})}_{\rm{test},k}$ and each label-wise centroid, respectively.
11:              $w_{jk}$ = label of the centroid for which the $2$-norm was smaller.
12:         end for
13:         Determine the label for every sample in $T^{(\ell,n,f)}_{\rm{test}}$ based on the one with which it is most associated in $w$, and receive a label sequence $L$.
14:     end for
15:     Compute MCC score between $L$ and $L^{\rm{(\rm{real})}}$ and store in $v$.
16:     Estimate $p$-value of the MCC score with hypothesis-testing routine and store in $u$.
17:end for
18:Apply FDR correction $u$ to receive an FDR-corrected array $u_{\rm{FDR}}$
19:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\rm{FDR}}$
Algorithm 4 Experiment: Supervised (Cosine Similarity)
1:Initiate MCC score array $v$ of length 100.
2:Initiate $p$-value array $u$ of length 100.
3:for $i=1,\ldots,100$ do
4:     Draw a subsample of $c\cdot\rm{min}\{m^{(\ell)}_{1},m^{(\ell)}_{2}\}$ text units for training ($T^{(\ell,n,f)}_{\rm{train}}$), the rest are for testing ($T^{(\ell,f)}_{\rm{test}}$).
5:     Compute two label-wise centroids of $T^{(\ell,f)}_{\rm{train}}$ according to $L^{\rm{(\rm{real})}}$.
6:     Initiate label-association array $w\in\mathbb{R}^{|T^{(\ell,f)}_{\rm{test}}|}$.
7:     for $j=1,\ldots,|T^{(\ell,f)}_{\rm{test}}|$ do
8:         $L_{j}=\rm{index}[\rm{max}(\rm{cos}(\rm{centroid}_{1},T^{(\ell,f)}_{\rm{test},j}),\rm{cos}(\rm{centroid}_{2},T^{(\ell,f)}_{\rm{test},j}))]$.
9:     end for
10:     Compute MCC score between $L$ and $L^{\rm{(\rm{real})}}$ and store in $v$.
11:     Estimate $p$-value of the MCC score with hypothesis-testing routine and store in $u$.
12:end for
13:Apply FDR correction $u$ to receive an FDR-corrected array $u_{\rm{FDR}}$
14:Chosen MCC score and $p$-value for the combination of $n$ and $\ell$, respectively: $\bar{v}$, $\bar{u}_{\rm{FDR}}$
5 Results
Here, we present the results of experiments where we (1) distinguish 50 pairs of texts composed by different authors and all (55) permutations of pairs of texts written by the same author in our corpus (§4.2), and (2) apply our hypothesis-testing method on all resulting classifications to determine whether it is based on predominantly sequentially correlated literary properties or not. We perform these experiments in four configurations, where in all cases, we split the texts into units of varying length (§4.4):
  •  $\bullet$
    Unsupervised Classification of Traditionally-Embedded Texts: We apply $2$-means algorithm (§4.5.1) to classify texts embedded according to either word $n$-grams or character $k$-mers of varying sizes and numbers of corpus-wise-weighed most-frequent features (§4.4.1).
  •  $\bullet$
    Supervised Classification of Traditionally-Embedded Texts: We apply the supervised GI framework (§4.5.2) to classify texts embedded according to either word $n$-grams or character $k$-mers of varying sizes and numbers of corpus-wise-weighed most-frequent features (§4.4.1).
  •  $\bullet$
    Unsupervised Classification of Neurally-Embedded Texts: We apply $2$-means algorithm (§4.5.1) to classify texts embedded using the pre-trained STAR model (§4.4.2).
  •  $\bullet$
    Supervised Classification of Neurally-Embedded Texts: We apply the supervised cosine similarity framework (§4.5.2) to classify texts embedded using the pre-trained STAR model (§4.4.2).
    In Figure 2, we demonstrate the output of one such experiment for the case of 50 word $n$-gram-embedded with $f$ = 300 pairs of texts composed by distinct authors, implementing the unsupervised approach. Results for $f$ = 100 and 200 yielded results nearly identical to those of $f$ = 300, such that we resolved not to provide a detailed account thereof as well. We summarize the important properties of our results for each experiment in the following manner:
  •  $\bullet$
    Binary True Positive Rate: The fraction of pairs of texts composed by different authors for which at least one combination of parameters yielded a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance, based on our hypothesis-testing framework.
  •  $\bullet$
    Total True Positive Rate: The fraction of the total combinations of parameters that yielded a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance of pairs of texts composed by different authors, based on our hypothesis-testing framework.
  •  $\bullet$
    Binary False Positive Rate: The fraction of pairs of texts composed by the same author for which at least one combination of parameters yielded a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance based on our hypothesis-testing framework.
  •  $\bullet$
    Total False Positive Rate: The fraction of the total combinations of parameters that yielded a classification of pairs of texts composed by the same author that is predominantly affected by non-sequentially correlated literary properties with high statistical significance, based on our hypothesis-testing framework.
  •  $\bullet$
    Binary True Negative: The fraction of pairs of texts composed by the same author for which no combination of parameters yielded a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance, based on our hypothesis-testing framework.
  •  $\bullet$
    Binary False Negative Rate: The fraction of pairs of texts composed by the different authors for which no combination of parameters yielded a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance, based on our hypothesis-testing framework.
    Figure 2: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of commingled texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using word $n$-grams with $f$ = 300. The x-axis is all combinations of features ($n$ and $l$), and the y-axis lists 50 chosen pairs of texts composed by distinct authors. Colored cells are combinations of parameters that yield a classification that is predominantly affected by non-sequentially correlated literary properties with high statistical significance. Significant cells are color-coded by, and list- the numerical value of the normalized MCC score. Blank cells represent classifications for which our hypothesis-testing framework yielded a $p$-value larger than 0.05.
    In Tables 2-4, we list all true/false positive rates for each experimental setup of traditionally- and neurally-embedded texts, respectively. Below, we discuss our results for the traditionally- and neurally-embedded texts, respectively.
    ($k$-means) Positive ($k$-means) Negative (GI) Positive (GI) Negative
    (Binary) True 100% 74.5% (85.4%) 100% 0% (0%)
    (Binary) False 25.5% (14.6%) 0% 100% (100%) 0%
    (Total) True 46.1% 91.7% (97.1%) 44.5% 84.2% (87.2%)
    (Total) False 8.3% (2.9%) 53.9% 15.8% (12.8%) 55.5%
    Table 2: Confusion matrices for word $n$-gram embedded texts for the unsupervised ($k$-means) and supervised (GI) classification approaches, with $f$ = 300 (see §4.4.1). True positives and false negatives refer to attempting to distinguish pairs of texts composed by different authors. In contrast, false positives and true negatives refer to attempting to distinguish pairs of texts composed by the same author. The bracketed rates indicate the results when overlooking false-positive classifications of texts of explicitly different genres, which we consider to be: [Lotr_1, Silmarillion], [Lotr_2, Silmarillion], [Lotr_3, Silmarillion], [Kipling_JungleBook, Kipling_Ballads], [Kipling_JungleBook2, Kipling_Ballads], [Narnia, Screwtape], [Caspian, Screwtape] (amounting to 16% of all pairs of texts composed by the same author in our corpus).
    ($k$-means) Positive ($k$-means) Negative (GI) Positive (GI) Negative
    (Binary) True 100% 70.9% (81.2%) 100% 63% (72.3%)
    (Binary) False 29.9% (18.8%) 0% 37% (27.7%) 0%
    (Total) True 65.1% 88.8% (93.9%) 65.1% 89.1% (94%)
    (Total) False 11.2% (6.1%) 34.9% 10.9% (6%) 34.9%
    Table 3: Confusion matrices for character $k$-mer embedded texts for the unsupervised ($k$-means) and supervised (GI) classification approaches, with $f$ = 300, similarly to Table 2.
    ($k$-means) Positive ($k$-means) Negative (cosine) Positive (cosine) Negative
    (Binary) True 100% 61.8% (70.8%) 92% 60% (68.8%)
    (Binary) False 38.2% (29.2%) 0% 30.9% (25%) 8%
    (Total) True 100% 65.8% (75.3%) 85.3% 64.2% (72.9%)
    (Total) False 34.2% (24.7%) 0% 35.8% (27.1%) 14.7%
    Table 4: Confusion matrices for STAR embedded texts for the unsupervised ($k$-means) and supervised (cosine) classification approaches, similarly to Table 2.
  • 5.1 Results: Traditionally-Embedded Texts
    An intriguing observation arises from the results, highlighting the remarkable true positive rates achieved by unsupervised methods, particularly notable in traditionally-embedded texts. This finding holds significance as unsupervised approaches, such as the $k$-means algorithm, operate without explicit optimization for authorship attribution tasks. Despite this, they exhibit robust performance in identifying texts by the same author, suggesting that these methods capture intrinsic textual similarities effectively. This success underscores the potential of unsupervised techniques in authorship attribution tasks, especially in scenarios where text representations are not specifically tailored for such purposes. As we demonstrate below, further exploration of these unsupervised methods could offer valuable insights into textual data’s underlying structures and patterns, enhancing our understanding of authorial styles and linguistic nuances.
    Another notable aspect of the results is the relatively low false positive rate observed in the unsupervised method, particularly evident in traditionally-embedded texts. Despite their inherent lack of optimization for authorship attribution, unsupervised methods exhibit the ability to minimize misclassifications of texts by different authors as authored by the same individual. This finding suggests that unsupervised approaches are adept at discerning subtle textual differences that distinguish authors, contributing to their effectiveness in authorship attribution tasks. That said, it is evidently a nontrivial task to discern between cases where false positive classifications are made based on genre-wise differences or not, as is demonstrated in our results.
    In contrast, the supervised GI method demonstrates a strong true positive identification capability, but its performance is marred by high false positive rates. This issue highlights a significant challenge faced by supervised classification approaches, where the classifier erroneously identifies texts by the same author as authored by different individuals. Such false positives can be attributed to various factors, including the complexity of linguistic features captured by the embedding techniques and the classifier’s susceptibility to common stylistic elements across an author’s works.
    In Figure 3, we apply the important features extraction method introduced in Yoffe et al. (2023) on the false positive classification between [Dickens_Copperfield, Dickens_OliverTwist], embedded using word $n$-gram size 2 with text unit length of 1000. This embedding yields an average MCC score of $\approx$90and an average $p$-value of $\approx 0.01$ – rendering it a classification that is not driven by sequential correlations with high statistical significance, despite the two texts having both been composed by Charles Dickens. In this case, as demonstrated in Figure 3, the distributions between first- and third-person speech features vary significantly, essentially a difference in genre stemming from the fact that David Copperfield is composed as a semi-auto-biography, leading to the false positive classification. Thus, by assessing the features yielding the classification – a considerably less challenging task undertaking in the case of traditional embeddings, such as word $n$-grams – interpretability analysis can provide valuable insight into the reason for false positive classifications and aid in tuning relevant feature subsets to avoid or detect them.
    Figure 3: Extracted important features for the classification of the texts pair [Dickens_Copperfield, Dickens_OliverTwist], embedded using word $n$-gram size 2 with text unit length of 1000.
    Our findings provide insight into the widespread effectiveness of various traditional embeddings in distinguishing between text units of varying lengths.
    In Figure 4, we illustrate the stylometric significance of different feature combinations (word $n$-grams/character $k$-mer sizes) and their average capacity to accurately differentiate texts composed by different authors.
    Our analysis reveals that for character $k$-mers, there is no significant difference in distinguishing power among sizes 2–6. However, all sizes demonstrate improved distinguishability when applied to smaller text units. Additionally, we observe that the decline in distinguishing power due to feature sparsity is not evident for $k=6$.
    Conversely, for word $n$-grams, we observe a distinct trend: Smaller text units exhibit better distinguishability, with the added factor of strong dependence on $n$-gram size. Particularly, there is a noticeable decrease in distinguishing power as $n$-gram size increases, likely attributed to feature sparsity effects.
    Figure 4: Distinguishing power of traditional parameter combinations: Left Panel: Character $k$-mers. Right Panel: Word $n$-grams.
    5.2 Results: Neurally-embedded Texts
    Here, an interesting result begs attention: in the unsupervised classification of pairs of text by different authors, neurally embedded texts achieve outstanding performance in binary true and total positive identification rates. In the supervised cosine case, however, both rates drop dramatically. Indeed, the expectation for neural embeddings coupled with supervised classification to outperform all other configurations proved false within our analysis’s scope and merits further exploration.
    In contrast to the traditionally-embedded texts scenario, the low false positive rates demonstrate that neural embeddings, such as STAR, can isolate the desired literary component (in this case – stylistic) and minimize the effect of undesirable influences such as those of thematic content. These results indicate the effectiveness of neural models tailored to the authorship-attribution task in minimizing misclassifications of texts by the same author as authored by different authors, even in supervised classification approaches.
    Comparing these results with the performance of supervised methods in traditionally-embedded texts, we observe a considerable disparity in false positive rates. While traditionally-embedded texts presented challenges for supervised methods, leading to relatively high false positive rates, neurally-embedded texts showcase the ability of supervised methods to achieve lower false positive rates. This discrepancy suggests that the inherent characteristics of neurally-embedded representations may offer advantages in mitigating false positives, thereby enhancing the reliability and accuracy of supervised authorship attribution systems. Further investigation into the underlying mechanisms driving these differences could yield valuable insights into the optimization and refinement of authorship attribution techniques across text embedding methodologies.
    One limitation inherent in the exploration of neurally-embedded texts is the challenge of interpretability, which restricts the depth of analysis, particularly regarding phenomena such as false positives. Unlike traditional embedding methods, which often provide explicit linguistic features that can be examined and manipulated to understand model behavior, neural embeddings operate more abstractly, making it challenging to dissect the underlying representations comprehensively. Consequently, while we could delve into false positives and their implications in the context of traditionally-embedded texts, a similar level of interpretability may not be achievable with neural embeddings. This limitation underscores the trade-off between the richness of representation offered by neural embeddings and the interpretability required for in-depth analysis, highlighting the need for innovative approaches to bridge this gap and unlock the full potential of neurally-embedded representations in authorship attribution tasks.
    6 Application to a Biblical Co-authored Text: The Priestly Source in the book of Exodus
    Scholars widely acknowledge that the current form of the Hebrew Bible is the product of numerous editorial processes, including additions and redactions. Consequently, it can be seen as a literary patchwork of co-authored documents, with sections that vary in genre, date, and origin. Many debates in this field have been ongoing for decades, if not centuries, attempting to delineate and distinguish atomic contributions, such as authors, schools of authors, or redactors, belonging to a specific historical, cultural, and religious context (e.g., Wellhausen, 1885; Gunkel, 1895; von Rad, 1973)).
    The Priestly source (often referred to as ”P” in scholarly discourse) is particularly notable within the study of the Hebrew Bible because it stands out as one of the most widely recognized and agreed-upon sources by biblical scholars (e.g., Haran, 1981; Campbell and O’Brien, 1993; Albertz, 2018). Unlike some other hypothetical sources in the Documentary Hypothesis (Baden, 2012), the Priestly source is distinguished by its distinct linguistic style, theological focus, and thematic consistency. P is characterized by its emphasis on ritual law, genealogies, and cultic practices, often displaying a formal, structured style that contrasts with the more narrative-driven or poetic portions of the Hebrew Bible. (e.g., Vink, 1969; Gertz, 2022). The agreement on the existence and characteristics of the Priestly source underscores its importance in the broader understanding of how the Hebrew Bible was composed and edited. It serves as a key piece in the larger puzzle of biblical composition, helping scholars trace the evolution of Israelite religion and identity through different historical periods.
    In previous work, we introduced an unsupervised pipeline employing $k$-means clustering to identify the combination of traditional features—such as text unit length and $n$-gram size—that best aligns with the hypothesized division of the book of Exodus into Priestly and non-Priestly components (Yoffe et al., 2023; Bühler et al., 2024), that is widely accepted by Biblical scholars (e.g., Smith, 1996). This approach achieved an optimal Matthews correlation coefficient (MCC) score of approximately 89%, with text units averaging 80 words in length and $n$-grams of size 2141414Detailed results concerning the optimal partitions, verses attributed to each cluster can be found at https://github.com/YoffeG/PnonP.
    Figure 5 presents the implementation of our hypothesis-testing approach to generate the null distribution of influences of sequentially correlated literary properties for the optimal unsupervised partition of the book of Exodus, which corresponds to the proposed distinction between Priestly and non-Priestly textual strata. The figure demonstrates that, although our method successfully generated label sequences with high MCC scores, the $p$-value associated with these results leads us to reject the hypothesis that the unsupervised partition of Exodus that best agrees with the hypothesized Priestly and non-Priestly strata is theme-based. Additionally, this null distribution can also be used to assess how far the partition departs from being affected by sequential correlations. If there are many high MCC scores within the null distribution, it suggests that such correlations are nontrivial and, subsequently, that the difference between the literary strata exhibits differences based on sequential correlations as well, albeit to a lesser extent.
    Figure 5: Null distribution that is comprised of 104 samples, of MCC scores for the unsupervised partitioning of the book of Exodus using $k$-means clustering and the hypothesized division of the book into a Priestly and non-Priestly textual components.
    7 Conclusions
    We propose a novel method for modeling sequential correlations within texts by encoding sequential dependencies of a reference label sequence into a correlated multivariate normal distribution. This allows us to stochastically draw label sequences that capture the sequential correlations of the original reference labels. This model forms the basis of a hypothesis-testing framework that assesses whether a given text classification is driven predominantly by sequential correlations.
    Through experiments on pairs of English prose texts—by both the same and different authors—we evaluate the method’s ability to distinguish classifications influenced by sequentially correlated literary properties, such as thematic content, typically expected in same-author texts within a similar genre, from those driven by other factors that are not sequentially correlated, such as stylistic differences in texts by different authors. We employ traditional word $n$-grams, character $k$-mers, and neural embeddings optimized for capturing authorial style, applying both supervised and unsupervised classification methods. Our approach effectively differentiates theme-driven classifications across all setups, underscoring its potential for enhancing authorship attribution by accurately modeling sequential correlations in texts and their effect on classification.
    Our findings highlight the effectiveness of unsupervised methods, particularly when applied to traditional embeddings, in authorship attribution tasks. Despite the absence of task-specific optimization, unsupervised methods show robust true positive rates and low false positives, reflecting their strength in capturing inherent textual patterns. While neural embeddings designed to isolate authorial style perform well in distinguishing texts by different authors, they can misclassify cases where texts by the same author differ by genre, resulting in some false positives. This suggests potential limitations in these embeddings’ sensitivity to nuanced thematic distinctions.
    Additionally, our analysis reveals that supervised classification methods, while effective in identifying true positives, are more prone to false positives, especially when complex or overlapping linguistic features are involved. This highlights the importance of interpretability in understanding the factors behind misclassifications. Our results also suggest that relying solely on feature distributions may be insufficient to separate style, genre, and theme contributions in classification outcomes.
    Further insights from traditional embeddings reveal that performance varies with text unit length. While character $k$-mers perform consistently across sizes, word $n$-grams perform better with smaller sizes, with diminishing distinguishing power as $n$ increases. Both embeddings are most effective within specific text length ranges, challenging the assumption that longer text units always yield better results. These findings inform feature selection strategies, enhancing model performance in stylometric tasks.
    We applied our method to the Book of Exodus, particularly to its Priestly and non-Priestly sections, widely considered distinct by biblical scholars. Our findings align with scholarly consensus, indicating that the classification difference between these components is not primarily sequentially correlated.
    In summary, our study demonstrates the potential of a data-centric hypothesis-testing approach to assess the influence of sequentially correlated literary properties in text classification. This method provides a refined understanding of how literary properties such as theme, style, and genre interact and influence classification. However, our results also highlight the challenge of fully disentangling these overlapping properties, particularly in supervised tasks. This work paves the way for future research on methods that can better separate these components, advancing the accuracy and interpretability of literary text analysis.
    Appendix A Text Pairs Tables
    Table A.1: Author-wise book groups from which multi-book samples of a single author are generated.
    Text (abbreviated) Group Number
    Potter_1 1
    Potter_2 1
    Potter_3 1
    Potter_4 1
    Potter_5 1
    Potter_6 1
    Potter_7 1
    Percy_1 2
    Percy_2 2
    Percy_3 2
    Percy_5 2
    Lotr_1 3
    Lotr_2 3
    Lotr_3 3
    Silmarillion 4
    Dickens_Copperfield 5
    Dickens_OliverTwist 5
    Dickens_HardTimes 5
    Dickens_TwoCities 5
    Dickens_Pickwick 5
    Holmes_1 6
    Holmes_2 6
    Holmes_3 6
    Kipling_JungleBook 7
    Kipling_JungleBook2 7
    Kipling_Ballads 8
    Narnia 9
    Caspian 9
    Screwtape 10
    Austen_Emma 11
    Austen_PP 11
    Austen_MansfieldPark 11
    spooky_EAP 12
    spooky_MWS 13
    spooky_HPL 14
    Table A.2: Randomly chosen pairs of texts composed by different authors that are evaluated in this work with their respective indices.
    Index Text 1 Text 2
    1 spooky_HPL Screwtape
    2 Lotr_3, Lotr_2, Lotr_1 Percy_5
    3 Austen_Emma, Austen_PP, Austen_MansfieldPark Screwtape
    4 Narnia, Caspian spooky_MWS
    5 Potter_4, Potter_2 Silmarillion
    6 spooky_MWS spooky_HPL
    7 spooky_HPL Potter_7, Potter_3, Potter_2
    8 spooky_MWS Lotr_3
    9 Dickens_TwoCities, Dickens_OliverTwist, Dickens_Copperfield Narnia, Caspian
    10 Austen_Emma Kipling_JungleBook
    11 spooky_MWS Narnia, Caspian
    12 Kipling_Ballads Lotr_1, Lotr_2, Lotr_3
    13 Kipling_Ballads Percy_1, Percy_3
    14 Potter_4, Potter_7, Potter_2 spooky_HPL
    15 Silmarillion Potter_7
    16 spooky_MWS spooky_HPL
    17 spooky_MWS Kipling_Ballads
    18 Lotr_2 Percy_5
    19 Narnia, Caspian spooky_HPL
    20 Potter_6, Potter_5, Potter_3 Holmes_1
    21 Austen_Emma, Austen_MansfieldPark, Austen_PP Dickens_TwoCities, Dickens_OliverTwist
    22 Percy_2, Percy_3, Percy_1 spooky_HPL
    23 Potter_1 Kipling_JungleBook2
    24 Holmes_1, Holmes_3, Holmes_2 Silmarillion
    25 spooky_EAP Kipling_Ballads
    26 spooky_EAP Potter_3, Potter_2
    27 Narnia, Caspian Kipling_Ballads
    28 spooky_HPL spooky_EAP
    29 Narnia, Caspian Potter_1, Potter_3
    30 Silmarillion spooky_HPL
    31 spooky_HPL Narnia, Caspian
    32 Kipling_JungleBook Screwtape
    33 Silmarillion spooky_MWS
    34 Lotr_2, Lotr_3 Kipling_JungleBook
    35 Lotr_2 Percy_5, Percy_2
    36 Kipling_Ballads Silmarillion
    37 Holmes_2, Holmes_3 spooky_HPL
    38 Screwtape Lotr_3, Lotr_1
    39 spooky_HPL Percy_3, Percy_2, Percy_5
    40 Austen_Emma, Austen_MansfieldPark spooky_HPL
    41 Lotr_1 Austen_PP, Austen_MansfieldPark
    42 Dickens_TwoCities Lotr_3, Lotr_1, Lotr_2
    43 spooky_HPL spooky_EAP
    44 Dickens_TwoCities, Dickens_HardTimes, Dickens_Pickwick Narnia, Caspian
    45 Screwtape spooky_MWS
    46 Narnia, Caspian Silmarillion
    47 Screwtape Narnia, Caspian
    48 Austen_PP Kipling_JungleBook2, Kipling_JungleBook
    49 Kipling_Ballads Narnia, Caspian
    50 Potter_2 Kipling_JungleBook2
    Table A.3: All permutations of same-author text pairs used in both experiments. Additionally, the fourth column lists our assumption regarding whether or not each pair of texts varies greatly in genre. When indecisive, we mark that pair with ”?”.
    Index Text 1 Text 2 Vary Genre-wise
    1 Potter_1 Potter_2 No
    2 Potter_1 Potter_3 No
    3 Potter_1 Potter_4 No
    4 Potter_1 Potter_5 No
    5 Potter_1 Potter_6 No
    6 Potter_1 Potter_7 No
    7 Potter_2 Potter_3 No
    8 Potter_2 Potter_4 No
    9 Potter_2 Potter_5 No
    10 Potter_2 Potter_6 No
    11 Potter_2 Potter_7 No
    12 Potter_3 Potter_4 No
    13 Potter_3 Potter_5 No
    14 Potter_3 Potter_6 No
    15 Potter_3 Potter_7 No
    16 Potter_4 Potter_5 No
    17 Potter_4 Potter_6 No
    18 Potter_4 Potter_7 No
    19 Potter_5 Potter_6 No
    20 Potter_5 Potter_7 No
    21 Potter_6 Potter_7 No
    22 Lotr_1 Lotr_2 No
    23 Lotr_1 Lotr_3 No
    24 Lotr_1 Silmarillion Yes
    25 Lotr_2 Lotr_3 No
    26 Lotr_2 Silmarillion Yes
    27 Lotr_3 Silmarillion Yes
    28 Percy_1 Percy_2 No
    29 Percy_1 Percy_3 No
    30 Percy_1 Percy_5 No
    31 Percy_2 Percy_3 No
    32 Percy_2 Percy_5 No
    33 Percy_3 Percy_5 No
    34 Dickens_Copperfield Dickens_OliverTwist ?
    35 Dickens_Copperfield Dickens_HardTimes ?
    36 Dickens_Copperfield Dickens_TwoCities ?
    37 Dickens_Copperfield Dickens_Pickwick ?
    38 Dickens_OliverTwist Dickens_HardTimes ?
    39 Dickens_OliverTwist Dickens_TwoCities ?
    40 Dickens_OliverTwist Dickens_Pickwick ?
    41 Dickens_HardTimes Dickens_TwoCities ?
    42 Dickens_HardTimes Dickens_Pickwick ?
    43 Dickens_TwoCities Dickens_Pickwick ?
    44 Holmes_1 Holmes_2 No
    45 Holmes_1 Holmes_3 No
    46 Holmes_2 Holmes_3 No
    47 Kipling_JungleBook Kipling_JungleBook2 No
    48 Kipling_JungleBook Kipling_Ballads Yes
    49 Kipling_JungleBook2 Kipling_Ballads Yes
    50 Narnia Caspian No
    51 Narnia Screwtape Yes
    52 Caspian Screwtape Yes
    53 Austen_Emma Austen_PP ?
    54 Austen_Emma Austen_MansfieldPark ?
    55 Austen_PP Austen_MansfieldPark ?
    Appendix B Traditional Embeddings Significance Maps
    B.1 Unsupervised Classification Results
    Figure B.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Figure B.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Figure B.3: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
    B.2 Supervised Classification Results
    Figure B.4: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
    Figure B.5: Significance map for the attempt to apply GI supervised classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Figure B.6: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using word $n$-grams with $f$ = 300, similarly to Figure 2.
    Figure B.7: Significance map for the attempt to apply GI supervised classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using character $k$-mers with $f$ = 300, similarly to Figure 2.
    Appendix C Neural Embeddings Significance Maps
    C.1 Unsupervised Classification Results
    Figure C.1: Significance map for the attempt to apply $2$-means classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
    Figure C.2: Significance map for the attempt to apply $2$-means classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.
    C.2 Supervised Classification Results
    Figure C.3: Significance map for the attempt to apply supervised cosine similarity classification to distinguish 50 pairs of texts composed by different authors (see Table LABEL:app_tab_unsupervised_texts_differentAuthors), embedded using STAR, similarly to Figure 2.
    Figure C.4: Significance map for the attempt to apply supervised cosine similarity classification to distinguish all permutations of pairs of texts composed by the same author (see Table LABEL:app_tab_unsupervised_texts), embedded using STAR, similarly to Figure 2.